# Follow the Mass - A Concordance Picture of Tidal Disruption Events

### v1

@@ 5.8 Detected event rates
Figure 3: Logarithmic contours of relative event rates $\partial^{2}{\dot{N}}_{\rm TDE}/\partial\log M_{\rm BH}\partial\log m_{\star}$ as found in flux-limited samples, assuming that the stellar mass-distribution has the same shape as a Salpeter IMF and that the intrinsic event rate per star per galaxy is independent of both the stellar and SMBH mass. Both $m_{\star}$ and $m_{{}_{\rm BH,6}}$ are plotted logarithmically.
@@
In the most naive model of the intrinsic TDE population, the probability of a TDE is independent of $m_{\star}$, so that the number of events is proportional to the IMF, but cut off at the maximum mass that has yet to leave the main sequence. The intrinsic population would therefore be dominated by low-mass stars. On the other hand, as we have just seen, we expect high-mass stars to yield more luminous events. The net effect on the observed population is to strongly limit the predominance of low-mass stars. Assuming a stellar mass distribution $dN/dm_{\star}\propto m_{\star}^{-2.35}$ (i.e., a young stellar population created with the Salpeter IMF) and that the intrinsic event rate is independent of $M_{\rm BH}$ (consistent with the predictions of Metzger & Stone (2016), Tucci & Volonteri (2017), and Polkas et al. (2023) for this range of black hole masses), the resulting relative rates (per $\log m_{\star}$ and $\log M_{\rm BH}$) are shown in Figure 3.
Despite the sharp decline with $M_{\star}$ of the Salpeter IMF, the luminosity bias places the highest event detection rate at $M_{\star}\sim 1-1.5M_{\odot}$ and $M_{\rm BH}\sim 0.5-2\times 10^{6}M_{\odot}$. The lowest rates, both a factor $\sim 3\times 10^{-3}\times$ the peak rate, are found at low stellar mass and high black hole mass ($M_{\star}\lesssim 0.5M_{\odot}$ and $M_{\rm BH}\gtrsim 3\times 10^{7}M_{\odot}$), and at high stellar mass and very low black hole mass: $M_{\star}\gtrsim 10M_{\odot}$ and $M_{\rm BH}\lesssim 2\times 10^{5}M_{\odot}$. Note that the rate for all $M_{\star}$ and $M_{\rm BH}\gtrsim 10^{7}M_{\odot}$ shown in Figure 3 is likely overestimated because we have not included the relativistic effects that magnify the fraction of events leading to direct capture (Kesden, 2012; Krolik et al., 2020).
@@ 6 Summary
In the past year, three global TDE simulations in which events with astrophysically-realistic parameters were studied ($M_{\star}\sim 1M_{\odot}$ and $M_{{}_{\rm BH}}\sim 10^{5}-10^{6}M_{\odot}$) have been published. Although they used different hydrodynamic algorithms (a fixed-grid intrinsically-conservative method, a moving-mesh intrinsically-conservative method, and SPH), in many respects their results are very consistent. In particular, all three find that after a few characteristic orbital times for the debris, the overwhelming majority of the debris mass—$\gtrsim 99\%$—is found at radii $\sim 3\times 10^{3}r_{\bf g}$. Thus, this fact is now a robust prediction of the basic physics of these events, subject to alteration only for extreme parameter choices (e.g., stellar pericenters $\lesssim 10r_{\bf g}$, for which strong apsidal precession would change the dynamics Krolik et al. (2020).).
@@
Because the location of the mass directly implies many physical quantities: orbital binding energy, orbital timescale, photospheric area, etc., the robust agreement on where the mass goes leads to a similarly robust agreement in their predictions of important observable properties. Although in all three simulations the specific parameters chosen were selected only for being within the plausible range and for computational convenience, the simulations’ predictions for the principal observables associated with the optical/UV flare lie squarely in the middle of the observed distribution (as compiled, e.g., by Hammerstein et al., 2023): total energy radiated a few ×$\times 10^{50}$ erg, peak optical/UV luminosity $\sim 1\times 10^{44}$ erg s-1, and temperature of a black-body component $\simeq 3\times 10^{4}$ K. The last prediction is particularly striking because it has almost no dependence on any of the parameters.
@@
In addition, the fact that nearly all the mass is deposited at such a large distance from the black hole directly implies that its subsequent accretion will stretch over a long period of time. It immediately follows that after the flare, the luminosity should decline over a period of several times the characteristic debris orbital timescale until it reaches rough constancy at a level $\sim 0.1L_{\rm peak}$. This prediction, too, is only weakly dependent on parameters and is nonetheless in excellent agreement with observations. In fact, because the total energy radiated on several year-timescales is generically comparable to the energy radiated during the flare (van Velzen et al., 2019a), observations demand a situation in which accretion of a large part of the debris mass takes place slowly.
@@
All three simulations predict similar luminosities because the underlying mechanism is also similar: shocks dissipating $\sim 10^{-3}$ of the debris’ rest-mass energy on a timescale $\sim t_{0}$. For this reason, all three likewise share a scaling relationship for how the luminosity depends on $M_{\star}$ and $M_{{}_{\rm BH}}$. Because the accessible volume for flux-limited surveys is $\propto L^{3/2}$, the rate at which events are detected in such surveys is $\propto L(M_{\star},M_{{}_{\rm BH}})^{3/2}\partial^{2}N_{\rm TDE}/\partial M_{\star}\partial M_{{}_{\rm BH}}$. Thus, the results of these simulations predict how much events with specific pairs of $M_{\star}$ and $M_{{}_{\rm BH}}$ are favored (or disfavored) for appearance in surveys, relative to their true rate. Interestingly, the rate enhancement is within a factor $\sim 3$ of its greatest amount for $0.6\lesssim M_{\star}/M_{\odot}\lesssim 6$ and $2\times 10^{5}\lesssim M_{{}_{\rm BH}}/M_{\odot}\lesssim 1\times 10^{7}$ if the intrinsic rate is proportional to the stellar IMF and independent of $M_{\rm BH}$. All the independent estimates of SMBH masses in the sample of Ryu et al. (2020b) are in this range, as are 28 of the 33 independently estimated SMBH masses in the Yao et al. (2023) compilation.
@@
We acknowledge support from National Science Foundation (NSF) grants AST-2009260 and PHY-2110339 (JHK). We also received support from the European Research Council Advanced Grant “MultiJets" and grant MP-SCMPS-00001470 from the Simons Foundation to the Simons Collaboration on Extreme Electrodynamics of Compact Sources - SCEECS (TP).
Appendix A Appendix - Relativistic Correction Factors
Two correction factors are important in estimating the depends of observational properties on the stellar and the black hole masses. The first factor determine the relation between the real tidal radius ${\cal R}_{T}$ and the order of magnitude estimate $r_{\rm t}$: ${\cal R}_{T}=\Psi(M_{\star},M_{{}_{\rm BH}})r_{\rm t}$ (Ryu et al., 2020c, d).
@@ Appendix A Appendix - Relativistic Correction Factors
$\Psi$ is a function of $m_{*}$, and possibly the star’s age and chemical composition, through the dependence of the star’s internal density profile on these factors. Taken as a function of $m_{\star}$ alone, it declines smoothly from $\approx 1.5$ for very small $m_{\star}$ to $\approx 0.43$ for all $m_{\star}\gtrsim 1.$. For a typical MS star encountering a black hole with $M_{\rm BH}=10^{6}M_{\odot}$, ${\cal R}_{T}\approx 25r_{\rm g}$ almost independent of the stellar mass (Ryu et al., 2020c, d; Law-Smith et al., 2020). This “physical tidal radius" is also a function of $M_{\rm BH}$ through the impact of general relativistic corrections; with $r_{\rm t}$ several tens of gravitational radii, they can be substantial (Ryu et al., 2020c, f): the portion of $\Psi$ dependent upon $M_{\rm BH}$ grows from $\approx 1.8$ for $m_{{}_{\rm BH,6}}\approx 10$ to $\approx 5$ for $m_{{}_{\rm BH,6}}=100$.
@@
Like the factor $\Psi$, $\Xi$ is a function of stellar mass, with additional possible dependence on stellar age and chemical composition. The function $\Xi(m_{\star})$ is almost opposite in behavior to $\Psi(m_{\star})$: $\Xi$ rises from an asymptote $\simeq 0.66$ for $m_{\star}\lesssim 0.3M_{\odot}$ to an asymptote at $\approx 1.8$ for $m_{\star}\gtrsim 1.3M_{\odot}$. $\Xi$ also depends upon $M_{\rm BH}$, particularly when $m_{{}_{\rm BH,6}}\gtrsim 1$, because of general relativistic effects (Ryu et al., 2020c, f): as a function of $M_{\rm BH}$, $\Xi$ falls from $\simeq 1.3$ in the Newtonian limit to $\approx 0.75$ when $m_{{}_{\rm BH,6}}=10$ and to $\approx 0.3$ at $m_{{}_{\rm BH,6}}=30$, beyond which the fitting formula no longer applies. Figure 4 illustrates how $\Xi$ is greatest for large $m_{\star}$ and small $m_{{}_{\rm BH,6}}$ and least for small $m_{\star}$ and large $m_{{}_{\rm BH,6}}$.

### v2

@@ 5.8 Detected event rates
Figure 3: Logarithmic contours of relative event rates $\partial^{2}{\dot{N}}_{\rm TDE}/\partial\log M_{\rm BH}\partial\log m_{\star}$ as found in flux-limited samples for a Salpeter mass function. The intrinsic event rate per star per galaxy is assumed to be independent of both the stellar and SMBH mass. Both $m_{\star}$ and $m_{{}_{\rm BH,6}}$ are plotted logarithmically ranging from 0.1 to 10.
@@
In the most naive model of the intrinsic TDE population, the probability of a TDE is independent of $m_{\star}$, so that the number of events is proportional to the stellar mass function. Because the lifetimes of stars with $m_{\star}\gtrsim 1$ are shorter than galaxy lifetimes, and higher mass stars yield more luminous events, the rate of events with high luminosity is very sensitive to the proportion of stars formed comparatively recently among the stars near enough the black hole to be disruption candidates. At the other extreme, the rate of events with low luminosity is sensitive to the slope of the mass function at small $m_{\star}$. Here we display in Figure 3 a simple example, a pure Salpeter mass function extending up to $10M_{\odot}$, and postpone more detailed discussions of event rates to another paper.
The shape of the detected events distribution in the $M_{\rm BH}\times M_{*}$ plane reflects two trade-offs. In rough terms, the dependence on $M_{*}$ balances the greater intrinsic number of low-mass stars against the higher luminosity of high-$M_{*}$ disruptions. The dependence on $M_{\rm BH}$ peaks at intermediate masses because events at lower black hole mass are limited to $\sim L_{E}\propto M_{\rm BH}$ and events at higher mass have lower peak luminosities because the timescale is longer. In this example using a Salpeter mass function, the luminosity bias places the highest event detection rate at $M_{\star}\sim 1-1.5M_{\odot}$ and $M_{\rm BH}\sim 0.5-2\times 10^{6}M_{\odot}$. The lowest rates are found at low stellar mass and high black hole mass ($M_{\star}\lesssim 0.5M_{\odot}$ and $M_{\rm BH}\gtrsim 5\times 10^{6}M_{\odot}$), and at high stellar mass and very low black hole mass: $M_{\star}\gtrsim 5M_{\odot}$ and $M_{\rm BH}\lesssim 2\times 10^{5}M_{\odot}$.
@@ 6 Summary
In the past year, three global TDE simulations in which events with astrophysically-realistic parameters were studied ($M_{\star}\sim 1M_{\odot}$ and $M_{{}_{\rm BH}}\sim 10^{5}-10^{6}M_{\odot}$) have been published. Although they used different hydrodynamic algorithms (a fixed-grid intrinsically-conservative method, a moving-mesh intrinsically-conservative method, and SPH), in many respects their results are very consistent. In particular, all three find that after a few characteristic orbital times for the debris, the overwhelming majority of the debris mass—$\gtrsim 99\%$—is found at radii $\sim 3\times 10^{3}r_{\rm g}$. Thus, this fact is now a robust prediction of the basic physics of these events, subject to alteration only for extreme parameter choices (e.g., stellar pericenters $\lesssim 10r_{\rm g}$, for which strong apsidal precession would change the dynamics Krolik et al. (2020).).
@@
Because the location of the mass directly implies many physical quantities: orbital binding energy, orbital timescale, photospheric area, etc., the robust agreement on where the mass goes leads to a similarly robust agreement in their predictions of important observable properties. Although in all three simulations the specific parameters chosen were selected only for being within the plausible range and for computational convenience, the simulations’ predictions for the principal observables associated with the optical/UV flare lie squarely in the middle of the observed distribution (as compiled, e.g., by Hammerstein et al., 2023a; Yao et al., 2023): total energy radiated a few ×$\times 10^{50}$ erg, peak optical/UV luminosity $\sim 1\times 10^{44}$ erg s-1, and temperature of a black-body component $\simeq 3\times 10^{4}$ K. This agreement with observed properties is especially striking because it is achieved entirely on the basis of physical calculations without any adjustment of parameters. Note, however, that our model may not be applicable to events in which $M_{{}_{\rm BH}}\gtrsim 10^{7}M_{\odot}$ because in this black hole mass range the critical pericenter for a full disruption is $<10r_{\rm g}$, a regime in which apsidal precession is strong enough to alter the flow pattern found in the three simulations. It is possible that the small number of events with $L_{\rm rad,peak}\gtrsim 10^{45}$ erg s-1 take place in this regime.
@@
In addition, the fact that nearly all the mass is deposited at such a large distance from the black hole directly implies that its subsequent accretion will stretch over a long period of time. It immediately follows that after the flare, the luminosity should decline over a period of several times the characteristic debris orbital timescale until it flattens out and maintains a roughly constant luminosity for many years. This prediction, too, is only weakly dependent on parameters and is nonetheless in excellent agreement with observations. In fact, because the total energy radiated on several year-timescales is generically comparable to the energy radiated during the flare (van Velzen et al., 2019a), observations demand a situation in which accretion of a large part of the debris mass takes place slowly.
@@
All three simulations predict similar luminosities because the underlying mechanism is also similar: shocks dissipating $\sim 10^{-3}$ of the debris’ rest-mass energy on a timescale $\sim t_{0}$. For this reason, all three likewise share a scaling relationship for how the luminosity depends on $M_{\star}$ and $M_{{}_{\rm BH}}$. Because the accessible volume for flux-limited surveys is $\propto L^{3/2}$, the rate at which events are detected in such surveys is $\propto L(M_{\star},M_{{}_{\rm BH}})^{3/2}\partial^{2}N_{\rm TDE}/\partial M_{\star}\partial M_{{}_{\rm BH}}$. Thus, the results of these simulations predict how much events with specific pairs of $M_{\star}$ and $M_{{}_{\rm BH}}$ are favored (or disfavored) for appearance in surveys, relative to their true rate; the highest luminosities are always strongly over-represented; events with 10×10$\times$10 × the typical luminosity are over-represented by a factor $\sim 30$. Interestingly, the rate at which events are found if the stellar mass function takes the Salpeter form and there has been relatively recent star formation near the galaxy center is within a factor $\sim 3$ of its greatest amount for $0.6\lesssim M_{\star}/M_{\odot}\lesssim 6$ and $2\times 10^{5}\lesssim M_{{}_{\rm BH}}/M_{\odot}\lesssim 1\times 10^{7}$. All the independent estimates of SMBH masses in the sample of Ryu et al. (2020b) are in this range, as are 28 of the 33 independently estimated SMBH masses in the Yao et al. (2023) compilation.
@@
We acknowledge support from the National Science Foundation (NSF) grants AST-2009260 and PHY-2110339 (JHK). We also received support from the European Research Council Advanced Grant “MultiJets" and grant MP-SCMPS-00001470 from the Simons Foundation to the Simons Collaboration on Extreme Electrodynamics of Compact Sources - SCEECS (TP).
Appendix A Appendix - Stellar Structure and Relativistic Correction Factors
Two correction factors are important in estimating the depends of observational properties on the stellar and the black hole masses. The first factor determine the relation between the real tidal radius ${\cal R}_{T}$ and the order of magnitude estimate $r_{\rm t}$: ${\cal R}_{T}=\Psi(M_{\star},M_{{}_{\rm BH}})r_{\rm t}$ (Ryu et al., 2020d, e).
@@ Appendix A Appendix - Stellar Structure and Relativistic Correction Factors
$\Psi$ is a function of $m_{*}$, and possibly the star’s age and chemical composition, through the dependence of the star’s internal density profile on these factors. Taken as a function of $m_{\star}$ alone, it declines smoothly from $\approx 1.5$ for very small $m_{\star}$ to $\approx 0.43$ for all $m_{\star}\gtrsim 1.$. For a typical MS star encountering a black hole with $M_{\rm BH}=10^{6}M_{\odot}$, ${\cal R}_{T}\approx 25r_{\rm g}$ almost independent of the stellar mass (Ryu et al., 2020d, e; Law-Smith et al., 2020). This “physical tidal radius" is also a function of $M_{\rm BH}$ through the impact of general relativistic corrections; with $r_{\rm t}$ several tens of gravitational radii, they can be substantial (Ryu et al., 2020d, c): the portion of $\Psi$ dependent upon $M_{\rm BH}$ grows from $\approx 1.8$ for $m_{{}_{\rm BH,6}}\approx 10$ to $\approx 5$ for $m_{{}_{\rm BH,6}}=100$.
@@
Like the factor $\Psi$, $\Xi$ is a function of stellar mass, with additional possible dependence on stellar age and chemical composition. The function $\Xi(m_{\star})$ is almost opposite in behavior to $\Psi(m_{\star})$: $\Xi$ rises from an asymptote $\simeq 0.66$ for $m_{\star}\lesssim 0.3M_{\odot}$ to an asymptote at $\approx 1.8$ for $m_{\star}\gtrsim 1.3M_{\odot}$. $\Xi$ also depends upon $M_{\rm BH}$, particularly when $m_{{}_{\rm BH,6}}\gtrsim 1$, because of general relativistic effects (Ryu et al., 2020d, c): as a function of $M_{\rm BH}$, $\Xi$ falls from $\simeq 1.3$ in the Newtonian limit to $\approx 0.75$ when $m_{{}_{\rm BH,6}}=10$ and to $\approx 0.3$ at $m_{{}_{\rm BH,6}}=30$, beyond which the fitting formula no longer applies. Figure 4 illustrates how $\Xi$ is greatest for large $m_{\star}$ and small $m_{{}_{\rm BH,6}}$ and least for small $m_{\star}$ and large $m_{{}_{\rm BH,6}}$.